OpenStudy (anonymous):

Calculate how much work you need to move a 280 N shed to a platform 5 m above

OpenStudy (anonymous):

We know that: \[Work = Force \times Distance\]\[W = F \times d\] Simply substitute known quantities and compute: \[W = F \times d\]\[W = 280N \times 5m\]\[W =1400J\]\[W = 1.4kJ\]